A mechanic can fix a car twice as fast as his helper. One day, the mechanic and helper worked together for 2 hours. Then the mechanic got called away, and the helper needed 1 more hour to finish the job. How long does it take the helper, working alone, to fix the car?
Answer by ikleyn(52778)

.
A mechanic can fix a car twice as fast as his helper.
One day, the mechanic and helper worked together for 2 hours.
Then the mechanic got called away, and the helper needed 1 more hour to finish the job.
How long does it take the helper, working alone, to fix the car?
~~~~~~~~~~~~~~~~

Let x be the rate of work of the helper; then the rate of work of the mechanic is 2x.


In 2 hours, the helper will do  2x parts of work; the mechanic will do 2*(2x) = 4x parts of work.


The total job equation is

    2x + 4x + x = 1,

or

        7x = 1,

which implies  x = 1/7.


It means that the helper needs 7 hours to do the job alone.    ANSWER

Solved.